# Towards Trapped-Ion Thermometry Using Cavity-Based EIT

### 09 Mar 2026 v3

@@ Abstract
We present a technique for measuring ion temperature using cavity-based electromagnetically induced transparency (EIT) applicable for cavity quantum electrodynamics (QED) systems in the strong coupling regime. This method enables efficient extraction of the ion’s phonon occupation number following sub-Doppler cooling close to the motional ground state. The proposed method relies on monitoring the cavity probe transmission while scanning the probe laser frequency to establish cavity-induced EIT using the control beam, significantly simplifying the measurement procedure. We establish a theoretical model that demonstrates the influence of the thermal state of the trapped ion vis-à-vis the EIT linewidth measured. We show through numerical simulations how the cavity-induced EIT transmission may be used as a thermometry tool to deduce the ion temperature as well as the motional state of the ion in the sub-Doppler cooling regime, even for systems that are in the weak coupling regime. The current method can only be used for operation in the resolved-sideband regime, where individual motional states can be selectively addressed for all relevant transitions either by selecting appropriate energy levels for the three-level system or by employing strong confinement with high secular frequencies ($\sim 10MHz$).
@@ I Introduction
Trapped ions offer a well-controlled platform for precision metrology, cavity quantum electrodynamics (QED), and quantum information processing due to their long coherence time and exceptional degree of experimental controllability [1, 2, 3, 4].
With successful demonstrations across a broad range of ion species and trap architectures, resolved sideband cooling has emerged as the standard technique for preparing trapped ions close to the motional ground state [5, 6, 7]. Achieving and confirming near-ground-state cooling is essential, since even a small residual motional excitation can degrade coherence and reduce the fidelity of quantum operations [7, 8]. For this reason, accurate thermometry after sideband cooling is crucial for describing cooling performance and validating subsequent quantum operations. After compensation of excess micromotion, it is possible to characterize the motion of a trapped ion in a radio-frequency Paul trap [7]. Ion temperature is directly determined by a discrete phonon occupation number that characterizes the motional state. Determining this phonon occupation provides a crucial benchmark for coherent quantum operations and enables a direct evaluation of the ions’ thermal state [1, 7, 9].
@@
A variety of techniques have been developed to deduce the motional occupation of trapped ions [10, 9, 11]. The most popular method is resolved sideband thermometry, which assumes a thermal motional distribution and uses the ratio of red to blue sideband transition strengths to estimate the mean phonon number [5, 12]. More generally, reconstruction techniques that do not strictly rely on thermal assumptions are made possible by extracting phonon distribution information from Rabi oscillations in carrier and sideband transitions [9, 10, 11]. These methods are well known and typically require precise internal-state preparation and the detection of projective states.
@@
Over the past few decades, following the seminal works of Kimble [13], Haroche [14] and Rempe [15], there has been growing interest in cavity quantum electrodynamics with trapped ions due to its potential for the realization of scalable quantum networks, enhanced light-matter coupling, and effective ion-photon interfaces for quantum information processing schemes and applications [16, 17, 18, 19, 20, 21, 22]. Interference-based phenomena such as cavity-assisted electromagnetic-induced transparency (EIT) are made possible by the presence of an optical cavity, which alters the interaction between the ion and the electromagnetic field. Since its discovery, EIT [23] has made it possible for a wide range of coherent optical control applications, such as slow light in atomic medium [24], quantum memory [25], laser cooling [26], precision spectroscopy, and sensing [16].
@@
Cavity-based EIT has been studied in a number of regimes to enable features such as narrowing of the cavity linewidth and enhanced optical nonlinearities [19, 20, 27]. A key feature of cavity-induced EIT is its sensitivity to decoherence mechanisms in the coupled ion-cavity system. Coupling to vibrational sidebands causes phonon-dependent dephasing of the effective $\Lambda$-type system in the presence of motional excitation, changing the coherence, and altering the EIT interference [28, 26]. The cavity-EIT transparency window is thus systematically expanded by thermal phonons. Thermally induced fluctuations alter the effective decoherence rates in the presence of motional excitation which affects the EIT transparency features’ width and shape. This dependence suggests that the cavity-EIT can serve as an indirect probe of the ion’s motional temperature.
@@
We subsequently extend the model to the multi-ion system by modifying the atom-cavity interaction Hamiltonian to include a summation over all ions coupled to the same cavity mode. In this formulation, each ion interacts with the cavity field identically, and the collective response of the ion ensemble naturally emerges from the full many-body master equation. This approach allows us to examine thermometry in regimes where the cavity response is improved by the collective light-matter interaction and the single-ion coupling strength to the intracavity field places the system in the weak coupling regime.
@@
The article is organized as follows: In section II, we present the theoretical model and Hamiltonian, including the treatment of ion motion and dissipative processes. In section III, numerical results for linewidth extraction and cavity transmission are shown. The section IV examines how phonons affect the cavity-EIT spectrum. The analysis of ion temperature and the role of thermalization dynamics are discussed in section V. The discussion then extends to multi-ion cavity-QED systems operating in the weak single-ion coupling regime in section VI, where we demonstrate that thermometry is made possible even in cases where individual coupling strengths are small through collective enhancement. section VII further investigates the applicability of the proposed method to systems with large excited state decay rates. We show conclusively that thermometry remains feasible under realistic experimental conditions. We present a realistic experimental realization roadmap in section VIII covering parameter regimes compatible with existing ion-cavity platform trap requirements and viable cavity configurations. Finally, a summary of the key results and concluding remarks are given in section section IX.
@@ II Theory and model
Cavity-based EIT systems have been extensively studied over the past two decades [19, 20, 29, 27, 30, 31, 32, 33]. These studies demonstrate that the cavity output can be effectively controlled by several key parameters, including the atom-cavity coupling strength ($g$), control field Rabi frequency ($\Omega_{c}$), and the cavity decay rate ($\kappa$). Among these parameters, the control field intensity, $\Omega_{c}$, represents an experimentally accessible degree of freedom, whereas the others are primarily defined by the cavity geometry, mirror reflectivity, and the choice of atomic energy levels. As a result, parameters such as $g$ and $\kappa$ cannot be easily tuned under typical experimental constraints, whereas $\Omega_{c}$ can be easily modified by tweaking the laser power. As the control field intensity $\Omega_{c}$ increases, the full width at half maximum (FWHM) of the EIT resonance also increases significantly, typically by $\approx 10~\mathrm{kHz}$. This dependence forms the basis of the proposed thermometry scheme.
For the purpose of ion thermometry, it is essential to include the vibrational levels associated with each relevant electronic energy level. Addressing these vibrational states requires operation in the resolved-sideband regime ($\gamma\ll\omega_{sec}$) [34, 1], where the splitting between adjacent vibrational levels, $\hbar\omega_{sec}$, exceeds the natural linewidth of the corresponding transition, $2\gamma$. Here, $\gamma$ denotes the spontaneous decay rate of the transition and $\omega_{sec}$ represents the secular frequency of the trapped ion.
Henceforth, a transition between two energy levels refers to a transition between their corresponding vibrational states. The allowed transitions can be classified into three main types: the carrier transition, the blue sideband (BSB) transition, and the red sideband (RSB) transition, corresponding to $\Delta n=0$ and $\Delta n=\pm 1$, respectively. When sideband transitions are addressed, the vibrational level occupancy, n, certainly influences the ion-field coupling strength. Consequently, the Rabi frequencies for the RSB and BSB transitions get modified as [1],
@@
Unlike the previous case, the Rabi frequency ($\Omega_{0}$) remains unchanged for the carrier transition. Thus, when the control laser is tuned to a sideband transition (BSB or RSB), the effect of the phonon occupancy, $n$, becomes evident in the EIT linewidth through its influence on the Rabi frequency, $\Omega_{c}$. In other words, variations in the EIT linewidth provide an estimate of changes in the ion’s vibrational occupancy. Motivated by this correlation, we define a systematic approach via numerical simulations for determining the ion temperature via phonon occupancy.
@@
where $\rho$ represents the density matrix of the ion-cavity system and $\kappa$ denotes the decay rate of the cavity field.
@@
The phonon states are inherent to damping due to fluctuating electric field noise present in real ion trap. To account for this scenario, we considered the trapped ion system such that it is coupled to a thermal bath. Fluctuating electric-field noise induces phonon exchange between the ion and the reservoir, driving the system toward thermal equilibrium [8].
@@
so close to the ground state ($\bar{n}=0$), $\dot{\bar{n}}(t)=\gamma_{b}n_{th}$. This gives an initial estimation of the heating rate of the ion. Equation 5, also shows that if a long time interval is taken, i.e., at equilibrium, $\bar{n}_{eq}=n_{th}$. So, determining the value of $n_{th}$ allows us to estimate the ion’s temperature owing to Equation 4.
@@ III Analysis of Transmission in Cavity-EIT
To study the system dynamics, we numerically solve the master equation (Equation 3) using the Hamiltonian given in Equation 2 for a single ion strongly coupled to the intracavity field. The probe-field detuning is varied across the cavity resonance while keeping the other system parameters fixed. The steady-state density matrix is obtained for each value of the probe detuning.
Figure 2: Simulated cavity-EIT transmission spectrum for $\kappa=0.4\,\mathrm{MHz}$, $g=3\kappa$, $\gamma_{eg}=\gamma_{eu}=\kappa$, $\langle n\rangle=1$, and $\gamma_{b}=0.25\kappa$. The observed linewidth ($0.55\,\mathrm{MHz}$) is narrower than the empty cavity linewidth $2\kappa=0.8\,\mathrm{MHz}$. The vacuum Rabi splitting peaks occur at $\pm\sqrt{g^{2}+\Omega_{c}^{2}}=\pm 1.56\,\mathrm{MHz}$.
From the steady-state solution, the photon population inside the cavity is extracted by evaluating the expectation value of the photon number operator, $\langle a^{\dagger}a\rangle$. The steady-state intra-cavity photon number, which is proportional to the output field leaking through the cavity mirrors, is then used to calculate the transmission spectrum. Figure 2) shows the resulting transmission profile. This spectrum exhibits the characteristic features of previously reported cavity-EIT systems [19, 36, 20, 27].
@@ V Determination of ion temperature and average phonon number
In addition, the average phonon occupancy, $\bar{n}$, is an important quantity for various applications, such as estimating heating rates, determining fidelity of ion trap-based quantum gates, indicating the Lamb-Dicke regime etc. Since the phonon occupancy is directly related to temperature (see Equation 4), $\bar{n}$ can also be estimated from the cavity-EIT linewidth.
Figure 4: (a) Cavity-EIT spectrum for different $n_{th}$ value (b) Ion temperature and average phonon occupancy ($\bar{n}$) as a function of cavity-EIT linewidth. The cavity decay rate, $\kappa=0.4MHz$, cavity coupling factor, $g=3\kappa$. The decay rates from the excited states are taken as $\gamma_{eg}=\gamma_{eu}=\kappa$. The phonon coupling rate is taken as $\gamma_{b}=0.25\kappa$.
We further perform simulations over a wide range of ion temperatures, from which the linewidth is calculated by fitting a Lorentzian function to the spectrum. The average phonon occupancy $\bar{n}$ is extracted by taking the expectation over phonon number operator ($\langle b^{\dagger}b\rangle$). In 4(b), the ion temperature and $\bar{n}$ are illustrated as functions of the cavity-EIT linewidth. The predicted temperature exhibits an asymptotic behavior at low linewidth values, indicating that at very low temperatures, variations in linewidth are minimal.
@@ VI Thermometry of a multi-ion cavity-QED system in the weak coupling regime
Figure 6: Temperature vs cavity-EIT linewidth for multi-ions. The EIT linewidth is scaled with respect to the empty cavity width. The typical single-ion cavity EIT linewidth for one temperature is shown in the inset with the set of different parameters. Except for temperature, the same parameters have been chosen for the simulations performed in this plot.
In the multi-atom/ion cavity system, the ion-cloud temperature may be estimated in a similar manner as outlined in the single-ion cavity EIT-based thermometry scheme as discussed in section V. We have shown the variation in temperature as a function of cavity-EIT linewidth in a fraction of empty cavity linewidth for different numbers of ions in Figure 6. The variation of the cavity EIT linewidth with temperature indicates that the linewidth is more sensitive to the temperature as one considers a higher number of atoms/ions coupled to the cavity mode. Thus, the proposed thermometry method is particularly well-suited for cavity-QED systems with low ion-cavity coupling strength, especially when multiple ions are coupled to the cavity mode.

### 26 Mar 2026 v4

@@ Abstract
We present a technique for measuring ion temperature using cavity-based electromagnetically induced transparency (EIT) applicable for cavity QED systems. This method enables efficient extraction of the ion’s phonon occupation number following sub-Doppler cooling close to the motional ground state. The proposed method requires operation in the resolved-sideband regime, where individual motional states can be selectively addressed for all relevant transitions either by selecting appropriate energy levels for the three-level system or by employing strong confinement with high secular frequencies ($\sim 10MHz$). It relies on monitoring the cavity probe transmission while scanning the probe laser frequency to establish cavity-induced EIT using a control beam, thereby significantly simplifying the measurement procedure. We establish a theoretical model that demonstrates the influence of the thermal state of the trapped ion vis-à-vis the EIT linewidth measured. We show through numerical simulations how the cavity-induced EIT transmission may be used as a thermometry tool to deduce the ion temperature as well as its motional state in the sub-Doppler cooling regime, even for systems that are in the weak coupling regime.
@@ I Introduction
Trapped ions offer a well-controlled platform for precision metrology, cavity quantum electrodynamics (QED), and quantum information processing owing to their long coherence time along with an exceptional degree of experimental controllability [1, 2, 3, 4].
With successful demonstrations across a wide range of ion species and trapping architectures, resolved sideband cooling has become the workhorse for preparing trapped ions close to the motional ground state [5, 6, 7]. Achieving and verifying near-ground-state cooling is of central importance, as even a small residual motional excitation can introduce decoherence and significantly reduce the fidelity of quantum operations [7, 8]. Consequently, precise thermometry following sideband cooling is essential for accurately characterizing the cooling performance and validating the initialization of the motional state prior to subsequent quantum manipulations.
Accurate knowledge of the ion temperature also plays a crucial role in several aspects of trapped-ion quantum technologies. In quantum information processing, precise thermometry helps mitigate motional decoherence and improves the fidelity of entangling gate operations [9, 10]. In ion-trap experiments, it enables the characterization of anomalous heating rates and provides insight into noise sources that limit trap performance [8]. Furthermore, in cavity-QED based quantum network protocols, controlling and diagnosing the motional state of ions is vital for optimizing ion–photon interfaces and ensuring efficient quantum state transfer between stationary and flying qubits [11, 12].
Post compensation of excess micromotion, the motion of an ion confined in a radio-frequency Paul trap can be described as a simple harmonic oscillator with discrete, equally spaced energy levels, which define the motional states [7]. Ion temperature is directly determined by a discrete phonon occupation number that characterizes the motional state. Determining this phonon occupation provides a crucial benchmark for coherent quantum operations and enables a direct evaluation of the ions’ thermal state [1, 7, 13].
@@
A variety of techniques have been developed to deduce the motional occupation of trapped ions [14, 13, 15]. The most popular method is resolved sideband thermometry, which assumes a thermal motional distribution and uses the ratio of red to blue sideband transition strengths to estimate the mean phonon occupation [5, 16]. More generally, reconstruction techniques that do not strictly rely on thermal assumptions are made possible by extracting phonon distribution information from Rabi oscillations in carrier and sideband transitions [13, 14, 15]. These methods are well known and typically require precise internal-state preparation and the detection of projective states.
@@
Over the past few decades, following the seminal works of Kimble [17], Haroche [18] and Rempe [19], there has been growing interest in cavity-QED with trapped ions due to its potential for the realization of scalable quantum networks, enhanced light-matter coupling, and effective ion-photon interfaces for quantum information processing schemes and applications [20, 21, 22, 23, 24, 25, 26]. Interference-based phenomena such as cavity-assisted electromagnetic-induced transparency (EIT) are made possible by the presence of an optical cavity, which alters the interaction between the ion and the electromagnetic field. Since its discovery, EIT [27] has made it possible for a wide range of coherent optical control applications, such as slow light in atomic medium [28], quantum memory [29], laser cooling [30], precision spectroscopy, and sensing [20].
@@
Cavity-based EIT has been studied in a number of regimes to enable features such as narrowing of the cavity linewidth and enhanced optical nonlinearities [23, 24, 31]. A key feature of cavity-induced EIT is its sensitivity to decoherence mechanisms in the coupled ion-cavity system. Coupling to vibrational sidebands causes phonon-dependent dephasing of the effective $\Lambda$-type system in the presence of motional excitation, changing the coherence and altering the EIT interference [32, 30]. The cavity-EIT transparency window is thus systematically expanded by thermal phonons. This sensitivity suggests that the cavity-EIT spectrum can provide a spectroscopic signature of the ion’s motional temperature.
Temperature measurements of trapped ions are often performed using resolved-sideband thermometry, which can obtain sensitivity close to the single-phonon level near the motional ground state, as discussed in [1]. Ion temperatures in the $10^{-1}-10^{2}$ mK range have also been measured with sub-mK resolution using dark-resonance thermometry, as shown in [33]. In trapped-ion systems imaging-based spatial thermometry typically offers precision at the millikelvin level [34]. Recently, cavity-based thermometry has been proposed, where the ion temperature is inferred from the sensitivity of the cavity emission to the ion’s spatial motion within the cavity mode [4]. In this method, the thermometry is based on a position-dependent cavity coupling resulting from the ions thermal motion.
In contrast, the thermometry scheme proposed in this article depends on the effective Rabi frequency in a cavity-EIT configuration being modified in a phonon-number dependent manner, which causes a detectable shift in the EIT linewidth. Aa an alternative thermometry method, particularly suited for cavity-QED experiments, our proposed method relies on measuring the cavity transmission spectrum with high frequency resolution. This enables even minute linewidth variations to correspond to discernible temperature changes.
@@
We subsequently extend the model to the multi-ion system by modifying the atom-cavity interaction Hamiltonian to include a summation over all ions coupled to the same cavity mode. In this formulation, each ion interacts with the cavity field identically, and the collective response of the ion ensemble naturally emerges from the full many-body master equation. This approach allows us to examine thermometry in regimes where the cavity response is strengthened by the collective light-matter interaction inspite of the single-ion coupling strength to the intracavity field placing the system in the weak coupling regime.
@@
The article is structured as follows: In section II, we demonstrate the theoretical model and Hamiltonian, including the treatment of ion motion and dissipative processes. In section III, numerical results for linewidth extraction and cavity transmission are shown. section IV examines how phonons affect the cavity-EIT spectrum. The analysis of ion temperature and the role of thermalization dynamics are discussed in section V. The discussion then extends to multi-ion cavity-QED systems operating in the weak single-ion coupling regime in section VI, where we demonstrate that thermometry is made possible even in cases where individual coupling strengths are small through collective enhancement. section VII further investigates the applicability of the proposed method to systems with large excited state decay rates. We show conclusively that thermometry remains feasible under realistic experimental conditions. In section VIII, we evaluate the projected sensitivity of the thermometry method. We present a realistic experimental realization roadmap in section IX covering parameter regimes compatible with existing ion-cavity platform trap requirements and viable cavity configurations. Finally, section X summarizes the main findings of this work and presents the concluding remarks.
@@ II Theory and model
Cavity-based EIT systems have been extensively studied over the past two decades [23, 24, 35, 31, 36, 37, 38, 39]. These studies demonstrate that the cavity output can be effectively controlled by several key parameters, including the control-field Rabi frequency ($\Omega_{c}$), the atom-cavity coupling strength ($g$), and the cavity decay rate ($\kappa$). Among these parameters, $\Omega_{c}$ represents an experimentally accessible degree of freedom, whereas the others are primarily defined by the cavity geometry, mirror reflectivity, and the choice of atomic energy levels. As a result, parameters such as $g$ and $\kappa$ cannot be easily tuned under typical experimental constraints, whereas $\Omega_{c}$ can be easily modified by tweaking the laser power. As $\Omega_{c}$ increases, the full width at half maximum (FWHM) of the EIT resonance also increases significantly, typically by $\approx 10~\mathrm{kHz}$. This dependence forms the basis of the proposed thermometry scheme.
For the purpose of ion thermometry, it is essential to include the vibrational levels associated with each relevant electronic energy level. Addressing these vibrational states requires operation in the resolved-sideband regime ($\gamma\ll\omega_{sec}$) [40, 1], where the splitting between adjacent vibrational levels, $\hbar\omega_{sec}$, exceeds the natural linewidth of the corresponding transition, $2\gamma$. Here, $\omega_{sec}$ represents the secular frequency of the trapped ion, and $\gamma$ denotes the spontaneous decay rate of the transition.
Henceforth, a transition between two energy levels refers to a transition between their corresponding vibrational states. The allowed transitions can be classified into three main types: the red sideband (RSB) transition, the blue sideband (BSB) transition, and the carrier transition, corresponding to $\Delta n=\mp 1$ and $\Delta n=0$, respectively. The vibrational level occupancy, $n$, directly influences the ion-field coupling strength when sideband transitions are addressed. Consequently, the Rabi frequencies for the RSB and BSB transitions get modified as [1],
@@
Unlike the previous case, the Rabi frequency ($\Omega_{0}$) remains unchanged for the carrier transition. Thus, when the control laser is tuned to the BSB or RSB transition, the effect of the phonon occupancy, $n$, becomes evident in the EIT linewidth through its influence on the Rabi frequency, $\Omega_{c}$. In other words, variations in the EIT linewidth provide an estimate of changes in the ion’s vibrational occupancy. Motivated by this correlation, we define a systematic approach via numerical simulations for determining the ion temperature via phonon occupancy.
@@
The phonon states are inherent to damping due to fluctuating electric field noise present in real ion trap. To account for this scenario, we considered the trapped ion system such that it is coupled to a thermal bath. Fluctuating electric-field noise induces phonon exchange between the ion and its surrounding reservoir, driving the system towards thermal equilibrium [8].
@@
so close to the ground state ($\bar{n}=0$), $\dot{\bar{n}}(t)=\gamma_{b}n_{th}$. This gives an initial estimation of the heating rate of the ion. Equation 5, also shows that if a long time interval is taken, i.e., at equilibrium, $\bar{n}_{eq}=n_{th}$. So, determining the value of $n_{th}$ provides an estimate of the ion temperature owing to Equation 4.
@@ III Analysis of Transmission in Cavity-EIT
To study the system dynamics, the master equation (Equation 3) has been solved numerically using the Hamiltonian given in Equation 2 for a single ion strongly coupled to the intracavity field. The probe laser frequency is scanned across the cavity resonance, keeping the other system parameters fixed. The steady-state density matrix is obtained for each value of the probe detuning.
Figure 2: Simulated cavity-EIT transmission spectrum for $\kappa=0.4\,\mathrm{MHz}$, $g=5\kappa$, $\gamma_{eg}=\gamma_{eu}=\kappa$, $\bar{n}=0.1$, and $\gamma_{b}=0.6\kappa$. The observed linewidth ($0.22\,\mathrm{MHz}$) is narrower than the empty cavity linewidth $2\kappa=0.8\,\mathrm{MHz}$. The vacuum Rabi splitting peaks occur at $\pm\sqrt{g^{2}+\Omega_{c}^{2}}=\pm 2.33\,\mathrm{MHz}$.
From the steady-state solution, the inter-cavity photon population is extracted by evaluating the expectation value of the photon number operator ($\langle a^{\dagger}a\rangle$). The steady-state intra-cavity photon number, which is proportional to the output field leaking through the cavity mirrors, is then used to calculate the transmission spectrum. In Figure 2 the resulting transmission profile is illustrated. This spectrum exhibits the characteristic features of previously reported cavity-EIT systems [23, 42, 24, 31].
@@ V Determination of ion temperature and average phonon number
In addition, the average phonon occupancy, $\bar{n}$, is an important quantity for various applications, such as estimating heating rates [11], determining the fidelity of ion trap-based quantum gates [48], indicating the Lamb-Dicke regime etc [1]. Since the phonon occupancy is directly related to temperature (see Equation 4), $\bar{n}$ can also be estimated from the cavity-EIT linewidth [49].
Figure 4: (a) Cavity-EIT spectrum for different $n_{th}$ value (b) Ion temperature and average phonon occupancy ($\bar{n}$) as a function of cavity-EIT linewidth. The cavity decay rate, $\kappa=0.4MHz$, cavity coupling factor, $g=5\kappa$. The decay rates from the excited states are taken as $\gamma_{eg}=\gamma_{eu}=\kappa$. The phonon coupling rate is taken as $\gamma_{b}=0.6\kappa$.
We further perform simulations over a wide range of ion temperatures, from which the linewidth is calculated by fitting a Lorentzian function to the spectrum. The average phonon occupancy $\bar{n}$ is extracted by taking the expectation over phonon number operator ($\langle b^{\dagger}b\rangle$). In 4(b), the ion temperature and $\bar{n}$ are illustrated as functions of the cavity-EIT linewidth. The predicted temperature exhibits an asymptotic behavior at low linewidth values, indicating that at high temperatures, variations in linewidth are minimal.
@@ VI Thermometry of a multi-ion cavity-QED system in the weak coupling regime
Figure 6: Temperature vs cavity-EIT linewidth for a multi-ion system. The EIT linewidth is scaled with respect to the empty cavity width. The typical single-ion cavity-EIT linewidth for a chosen temperature is shown in the inset with other system parameters. Except for temperature, the same parameters have been chosen for the simulations performed in this plot.
In the multi-atom/ion cavity system, the ion-cloud temperature may be estimated in a similar manner as outlined in the single-ion cavity EIT-based thermometry scheme as discussed in section V. in Figure 6, we have shown the variation in temperature as a function of cavity-EIT linewidth in a fraction of empty cavity linewidth for different numbers of ions. The variation of the cavity EIT linewidth with temperature indicates that the linewidth is more sensitive to the temperature as one considers a higher number of atoms/ions coupled to the cavity mode. Thus, the proposed thermometry method is particularly well-suited for cavity-QED systems with low ion-cavity coupling strength, especially when multiple ions are coupled to the cavity mode.